Shift report 20000818 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Aug 18 2000 - 08:07:25 EDT

  • Next message: http account: "Shift report 20000818 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Aug 18 08:07:25 2000, Shift: 00:00-08:00
    Supervisor: FR, with: EJ
    Log book#4 pages 115 - 115
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was cogged from 8pm (previous shift) until 5.40am.
       The trigger 1 rate decreased from about 60 at the beginning
       of the store to about 10 at the end. 
    
    2. Run 2336 : MRS at 90deg., 4KGauss, MTPC1 anode  voltage
        at -1200V. It contains about 220K collisions.
    
    3. T3 cannot hold any voltage, set off.
        No other detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Aug 18 2000 - 08:10:52 EDT